# A Classification Theorem for Steady Euler Flows

### v1

@@ 4.1 The Unextendable Case
Lemma 4.10 shows there exists $\lambda_{0}>0$ with $h_{\lambda_{0}}\geq 0$ in $\breve{\Omega}_{\lambda_{0}}$ and at least one of the internal boundary tangency and boundary tangency happens for $\breve{\Omega}_{\lambda_{0}}$. In the case where internal tangency occurs at $x^{0}\in\partial\Omega_{\Gamma}$, by Lemma 4.11, $h_{\lambda_{0}}$ is a non-negative solution of the equation $\Delta h_{\lambda_{0}}=c(x)h_{\lambda_{0}}$ in $\breve{\Omega}_{\lambda_{0}}$, with $c^{-}(x)\leq\frac{M}{dist(x,\partial\breve{\Omega}_{\lambda_{0}})}$. Moreover, we have $\nabla h_{\lambda_{0}}(x^{0})=\nabla\psi(x^{0})-\nabla\psi_{\lambda_{0}}(x^{0})=0$. Now we apply Proposition A.1 for a sufficiently small ball $B\subset\breve{\Omega}_{\lambda_{0}}$ that hits $\breve{\Omega}_{\lambda_{0}}$ internal-tangentially at $x^{0}$, and conclude $h_{\lambda_{0}}=0$. Hence $\psi$ and correspondingly $\Omega_{\Gamma}$ is even symmetric in the direction $(1,0)$.
Now we consider the case where the boundary of tangency happens at $x^{0}\in H_{\lambda_{0}}\cap\partial\Omega_{\Gamma}$. First, similar to above, we have $\nabla h_{\lambda_{0}}(x^{0})=0$. In addition, $\partial\breve{\Omega}_{\lambda_{0}}$ forms a right angle at $x^{0}$, and the normal direction of $\partial\Omega_{\Gamma}$ at $x^{0}$ is parallel to direction $(0,1)$ In particular, we have $\nabla^{2}h_{\lambda_{0}}(x^{0})=0$. Now analogously as in the case of the internal tangency, by applying Proposition A.2 in the appendix to a small half ball contained in $\breve{\Omega}_{\lambda_{0}}$ that hits $\Omega_{\Gamma}$ boundary-tangentially, we have $h_{\lambda_{0}}=0$. Hence $\psi$ and correspondingly $\Omega_{\Gamma}$ is even symmetric in the direction $(1,0)$. ∎
4.2 The Extendable Case
Now we assume that $F$ is analytic at $s=0.$ Using the argument in the proof of Proposition 2.2, we see that we can extend the definition of $F$ so that $\psi$ satisfies
$\Delta\psi=F(\psi)$
on an extended domain $\tilde{\Omega}_{\Gamma}$, whose boundary is again either an analytic curve of points of degree 2, or a curve of points of degree 3 or higher. By the Jordan curve theorem, $\Gamma$ is either in the interior of the $\partial\Omega_{\Gamma}$ or in the exterior. In either case, since we are assuming that $F$ cannot be extended to all of $\Omega$, we will hit a curve of critical points along which $F$ is not analytic. In that case, we take the smallest curve inscribed in it with that property and we apply either Proposition 3.7 or 4.3 depending on the degree of vanishing of $\nabla\psi$ on the boundary. The proof of Theorem 1.1 is now complete.
4.3 Counterexample
@@
It is important to note that the rigidity in Theorem 1.1 is optimal in terms of regularity in bounded domains. First, the results of [18] show that Theorem 1.1 does not hold for flows that have finite smoothness even for the weaker assumption of non-locally-radial flows. Secondly, Theorem 1.1 does not hold in the case where the compact domain $\Omega$ is not simply connected, as the following result shows.
Theorem 4.12.
@@
Fix an analytic Jordan curve $\gamma\subset\mathbb{R}^{2}$. Then there exists a tubular neighborhood $\Omega$ of $\gamma$ and a stream function $\psi\in C^{\omega}(\bar{\Omega})$ constant on $\partial\Omega$ solving
$\begin{cases}\Delta\psi=2+\psi^{\frac{5}{2}}\text{ in }\Omega^{interior}\\ \Delta\psi=2-\psi^{\frac{5}{2}}\text{ in }\Omega^{exterior}\end{cases},\text{ and }\psi=\frac{\partial\psi}{\partial n}=0\text{ on }\gamma.$
@@
Thus, any non-circular $\gamma$ gives rise to a steady Euler flow $\psi$ that is analytic in a neighborhood of $\gamma$ that is neither radial nor a solution of a semilinear elliptic equation of the form (1.4). The proof of Theorem 4.12 is inspired by the classical proof the Cauchy-Kovalevskaya Theorem and a careful application of the so-called majorization method to a well chosen majorization system. A detailed proof is given in the thesis of the second named author [26].
5 Conclusion and Future Directions
@@
We gather here a few interesting directions and problems that can be considered, in view of Theorem 1.1 and its proof. We have already mentioned that replacing the analyticity assumption with some other less stringent assumption would be desirable. We mention some other interesting problems below.
5.1 Local and Global Questions
@@
$\mathcal{L}=\Delta-F_{*}^{\prime}(\psi_{*})$
play an important role in the local structure of steady states (see [9, 10, 11, 12, 34]). One consequence of Theorem 1.1 and the proof is that we can always define $\mathcal{L}$ as a (unbounded) self-adjoint operator on $L^{2}$. This gives hope to have a general local understanding of the manifold of steady states, and then possibly a global one.
Most of the previous results on the local structure of steady states have assumed the invertibility of $\mathcal{L}$. It is a very interesting problem to study the general case where $\mathcal{L}$ has a kernel. From the classical Fredholm theory, we know the kernel is finite dimensional. When the kernel of $\mathcal{L}$ is one-dimensional, it is still possible to give a quite clear picture of the local structure of steady states (though this has not yet been spelled out in the literature). Once the kernel of $\mathcal{L}$ is two-dimensional or higher, the local structure of steady states may be quite complicated [11]. It is a very interesting problem to study this in general and this is related to the (im)possibility of the existence of isolated states [14]. If we had a good understanding of the local structure of steady states in the general case, we could hope to patch together this information to get more global information about the structure of the set of analytic Euler steady states. It is possible that some of these issues could be simpler to study on generic domains [42]. For example, on generic domains it may even be possible to establish some kind of global correspondence between $\psi$ and $F$ in the analytic class [9].
In general, it would be very interesting to have results related to the connectedness, path-connectedness, and smoothness of the manifold of steady states [12].
5.2 Rigidity questions on other domains
A first obvious question is whether similar results can be established in different domains. We have already shown that it does not hold on some annular domains, though it is not clear whether a modification holds on $\mathbb{T}\times[0,1],$ for example. A very interesting problem is to determine whether a version of Theorem 1.1 can hold on $\mathbb{T}^{2}.$ In that case, the question would be whether all analytic steady states are necessarily either a shear flow or a solution to a semilinear elliptic equation.
@@ Acknowledgements
T.M.E. acknowledges partial funding from the NSF DMS-2043024, the Alfred P. Sloan foundation, and the Simons Foundation. He also acknowledges helpful conversations with T. Drivas and A. Kiselev. Y.H. acknowledges partial funding from NSF DMS-2043024. A. R. S. is supported by the ERC/UKRI advanced grant SWAT. The research of C. X. is partially supported by NSFC grant 12250710674, and Program of Shanghai Academic Research Leader 22XD1421400.
Appendix A Hopf Lemmas for elliptic equations with singular coefficents
In this section, we list some known facts about the maximum principle for the solutions to the elliptic equation $-\Delta h+c(x)h=0$ in a Lipschitz domain $\Omega$. Here $c(x)$ could be unbounded, but it satisfies the following uniform upperbound $c^{-}(x)\leq\frac{M}{dist(x,\partial\Omega)}$. The results we present here are crucial in our proof of Proposition 4.3.
@@ Appendix A Hopf Lemmas for elliptic equations with singular coefficents
$-\Delta h+c(x)h=0\text{ with $c^{-}(x)\leq\frac{M}{dist(x,0)}$.}$ (A.2)
Now if we further assume $\nabla h(0)=h(0)=0$, there exists a $r_{0}>0$ depending on $M$, such that if $r<r_{0}(M)$, we have either $h=0$ or $\frac{\partial^{2}h}{\partial\eta^{2}}(0)>0$, for any direction $\eta=(\eta_{1},\eta_{2})$ satisfying $\eta_{1}>0$, $\eta_{2}<0$ in $B_{r}^{+}$.

### v2

@@ 4.1. The Unextendable Case
Without loss of generality, we assume $c=0$. Since $F$ is not analytic at 0, $\partial U$ consists of critical points of $\psi$. By Proposition 2.3, we can find an innermost loop $\Gamma_{0}$ of critical points of $\psi$ where $F$ loses analyticity (i.e., $F$ is analytic at the values of $\psi$ in every curve of critical points contained in the open region enclosed by $\Gamma_{0}$). Now $\psi$ must be single-signed in the region $\Omega_{\Gamma_{0}}$ which is enclosed by $\Gamma_{0}$. Otherwise, there exist two points $x^{0},x^{1}\in\Omega_{\Gamma_{0}}$ such that $\psi(x^{0})<0$ and $\psi(x^{1})>0$. By the continuity of $\psi$, there is a zero point of $\psi$ in every continuous curve that connects $x^{0}$ and $x^{1}$ and is contained in $\Omega_{\Gamma_{0}}$. In particular, $\{\psi=0\}$ has infinitely many points in $\Omega_{\Gamma_{0}}$ and there exists a curve $\Gamma_{1}$ consisting of zeros of $\psi$. Since $F$ is not analytic at 0, $\Gamma_{1}$ must be a curve of critical points. This contradicts the definition of $\Gamma_{0}$. Analogously as in Section 3.2 and Section 4.1, applying Proposition 3.6 or Proposition 4.3 depending on the degree of vanishing of $\psi$ in $\Gamma_{0}$, $\psi$ must be a radially symmetric function. ∎
To prove Theorem 1.1, we need the following important lemma.
Lemma 4.10.
Assume that $U$ is a simply connected bounded open set, $\psi\in C^{\omega}(\bar{U})$ is constant on $\partial U$, and $\{\psi,\Delta\psi\}=0$ in $U$. If $\psi$ is neither radially symmetric nor realizable in $U$, then there exists a simply connected domain $\tilde{U}\varsubsetneq U$ such that $\nabla\psi=0$ on $\partial\tilde{U}$ and $\psi$ is not realizable in $\tilde{U}$.
Proof.
Suppose that there exists an $x\in U$ such that $\nabla\psi(x)\neq 0$, otherwise, $\psi$ is constant and thus realizable in $U$. Then, by Lemma 1.3, $\psi$ is realizable by some $F$ in $B_{\epsilon}(x)$ for some small $\epsilon.$ Now let $U_{0}$ be the largest connected set containing $B_{\epsilon}(x)$, in which $\psi$ is realizable. By analyticity of $\psi$, $\psi$ is relizable by $F$ in $U_{0}$. Since $\psi$ is not realizable in $U$, $U_{0}$ must be a proper subset of $U$.
We claim that $U_{0}$ cannot be simply connected. Indeed, if $U_{0}$ were simply connected, then $\partial U_{0}$ is a loop of critical points of $\psi$ and that $F$ is not analytic at $c=\psi(\partial{U}_{0})$. It follows from Proposition 4.9 that $\psi$ must be radially symmetric. This leads to a contradiction.
We now take $\tilde{U}$ to be a simply-connected component of the set $\Omega\setminus U_{0}$. $\tilde{U}$ would be a simply connected domain, and the boundary $\tilde{\Gamma}$ is a piesewise analytic curve. Moreover, assume $\psi$ obtain value $\tilde{c}$ at $\tilde{\Gamma}$, then the $F$ ($F$ is previously defined in the range of $\psi$ in $\overline{U_{0}}$ and we have $\Delta\psi=F(\psi)$ in $\overline{U_{0}}$.) is not analytic at the value $\tilde{c}$. Now, suppose $\psi$ were realizable in $\tilde{U}$, say by $G$. We claim that $G$ is not analytic at $\tilde{c}$. Otherwise, let $\tilde{G}$ be the analytic extension of $G$. As $\Delta\psi-\tilde{G}(\psi)$ is an analytic function in a tubular neighborhood $\mathcal{N}$ of $\tilde{\Gamma}$ and $\Delta\psi-\tilde{G}(\psi)=0$ in $\tilde{U}$, we have $\Delta\psi-\tilde{G}(\psi)=0$ in $\mathcal{N}$. This, in particular, implies that $F$ is analytic at $\tilde{c}$, which leads to a contradiction. Since $G$ is not analytic at $\tilde{c}$, we could apply Proposition 4.9 to $\tilde{U}$ and conclude that $\psi$ is radially symmetric. Therefore, $\psi$ is not realizable in $\tilde{U}$. Clearly, by definition, $\tilde{U}\varsubsetneq U$. This completes the proof of the lemma. ∎
Based on Lemma 4.10, we give the proof of Theorem 1.1.
Proof of Theorem 1.1.
Suppose $\psi$ is neither radially symmetric nor realizable in $\Omega_{0}:=\Omega$, applying Lemma 4.10 iteratively, we obtain a nested family of simply connected domains $\{\Omega_{n}\}_{n=1}^{\infty}$ such that for $n\geq 1$, $\Omega_{n}\subsetneq\Omega_{n-1}$, $\psi$ is not realizable in $\Omega_{n}$, and $\nabla\psi=0$ on $\partial\Omega_{n}$. This, in particular, implies that there is an infinite family of nested loops of critical points for the analytic function $\psi$. It follows from Proposition 2.3 that $\psi$ is a constant function. This contradiction implies that $\psi$ must be radially symmetric or realizable in $\Omega$. ∎
4.3. Counterexample
@@
It is important to note that the rigidity in Theorem 1.1 is optimal in terms of regularity in bounded domains. First, the results of [23] show that Theorem 1.1 does not hold for flows that have finite smoothness, even for the weaker assumption of non-locally-radial flows. Secondly, Theorem 1.1 does not hold in the case where the compact domain $\Omega$ is not simply connected, as the following result shows.
Proposition 4.11.
@@
Fix an analytic Jordan curve $\gamma\subset\mathbb{R}^{2}$. Then there exist a tubular neighborhood $\Omega:=\Omega^{int}\cup\gamma\cup\Omega^{ext}$ of $\gamma$ and a stream function $\psi\in C^{\omega}(\bar{\Omega})$, which is constant on $\partial\Omega$ and solves the following problem.
[…]
@@
Thus, any non-circular $\gamma$ gives rise to a steady Euler flow $\psi$, that is analytic in a neighborhood of $\gamma$, and that is neither radial nor a solution of a semilinear elliptic equation of the form (1.4).
In order to prove Proposition 4.11, we construct an analytic function $\psi$ in a tubular neighborhood $\overline{\Omega^{int}\cup\Omega^{ext}}$ of $\gamma$ satisfying
$\Delta\psi=2+\psi^{\frac{5}{2}}\quad\text{in }\Omega^{int}.$
The construction of such an analytic function is inspired by the classical proof of the Cauchy–Kovalevskaya Theorem, where we apply the so-called majorization method to a carefully chosen majorizing system, so that the existence of the analytic solution is established.
Once such an analytic function $\psi$ is obtained, we can verify that $\{\psi,\Delta\psi\}=0$ in $\Omega^{ext}$. Analogous to the proof of Lemma 4.4, we can then prove
$\Delta\psi=2-\psi^{\frac{5}{2}}\quad\text{in }\Omega^{ext}.$
A detailed proof can be found in Sections 2.4.3–2.4.5 of the thesis of the second-named author [31].
5. Conclusion and Future Directions
@@
We gather here a few interesting directions and problems that can be considered in view of Theorem 1.1 and its proof. We have already mentioned that replacing the analyticity assumption with some other less stringent assumption would be desirable. We mention some other interesting problems below.
5.1. Local and Global Questions
@@
$\mathcal{L}=\Delta-F^{\prime}(\psi_{*})$
play an important role in the local structure of steady states (see [12, 13, 15, 16, 36]). One consequence of Theorem 1.1 and the proof is that we can always define $\mathcal{L}$ as a (unbounded) self-adjoint operator on $L^{2}$. This gives hope for a general local understanding of the manifold of steady states and then possibly a global one.
Most of the previous results on the local structure of steady states have assumed the invertibility of $\mathcal{L}$. It is an interesting problem to study the general case where $\mathcal{L}$ has a kernel. From the classical Fredholm theory, we know that the kernel is finite-dimensional. Once $\mathcal{L}$ has a kernel, the local structure of steady states may be quite complicated [15]. It is an interesting problem to study this in general, and this is related to the (im)possibility of the existence of isolated states [20]. If we had a good understanding of the local structure of steady states in the general case, we could hope to patch together this information to get more global information about the structure of the set of analytic Euler steady states. Some of these issues could be simpler to study in generic domains [49]. In general, it would be interesting to have results related to the connectedness, path-connectedness, and smoothness of the manifold of steady states (see [16, 17]).
5.2. Rigidity Questions on Other Domains
A first obvious question is whether similar results can be established in different domains. We have already shown that it does not hold in some annular domains, although it is not clear whether a modification holds on $\mathbb{T}\times[0,1],$ for example. An interesting problem is to determine whether a version of Theorem 1.1 can hold on $\mathbb{T}^{2}.$ In that case, the question would be whether all analytic steady states are necessarily either a shear flow or a solution to a semilinear elliptic equation.
@@ Acknowledgements
T. M. E. acknowledges partial funding from the NSF DMS-2043024, the Alfred P. Sloan foundation, and the Simons Foundation. He also acknowledges helpful conversations with T. Drivas and A. Kiselev. Y. H. acknowledges partial funding from NSF DMS-2043024. A. R. S. is supported by the ERC/UKRI advanced grant SWAT. The research of C. X. is partially supported by NSFC grant 12250710674, and Program of Shanghai Academic Research Leader 22XD1421400. The authors thank the referees for many helpful comments that significantly improved the presentation of this paper.
Appendix A Hopf Lemmas for Elliptic Equations with Singular Coefficients
In this section, we list some known facts about the maximum principle for solutions to the elliptic equation $\Delta h+c(x)h=0$ in a Lipschitz domain $\Omega$. Here $c(x)$ could be unbounded but satisfies the following uniform upper bound $c^{-}(x)\leq\frac{M}{dist(x,\partial\Omega)}$. The results we present here are crucial in our proof of Proposition 4.3.
@@ Appendix A Hopf Lemmas for Elliptic Equations with Singular Coefficients
(A.2) $\Delta h+c(x)h=0\text{ with $c^{-}(x)\leq\frac{M}{dist(x,0)}$.}$
Now, if we further assume $\nabla h(0)=h(0)=0$, there exists a $r_{0}>0$ depending on $M$, such that if $r<r_{0}(M)$, we have either $h=0$ or $\frac{\partial^{2}h}{\partial\eta^{2}}(0)>0$, for any direction $\eta=(\eta_{1},\eta_{2})$ satisfying $\eta_{1}>0$, $\eta_{2}<0$ in $B_{r}^{+}$.
Appendix B On the Rigidity of Steady States with Single-Signed Vorticity
We now proceed to make precise statements in Remark 1.5.
Proposition B.1.
Let $\Omega$ be a simply connected, bounded, and convex domain, and $\psi$ be the stream function of a steady Euler flow with non-positive vorticity in $\Omega$. If we further assume that $\Omega$ is even symmetric in the direction $e:=(1,0)$, and $\psi$ is analytic, then $\psi$ is even symmetric in the direction $e$.
The full proof of this theorem follows inverbatim from the proof of Theorem 1.1. We explain the key ideas here. In the case where $\psi$ is nonradial, by Theorem 1.1, there exists an $F$ such that $\Delta\psi=F(\psi)$ in $\Omega$. Without loss of generality, we assume $\psi=0$ in $\partial\Omega$. As $\Delta\psi$ is non-positive, unless $\psi=0$, we have $\psi>0$ in $\Omega$ and $F$ is analytic near 0. Without loss of generality, $max_{\Omega}\psi=1$, then $F$ is analytic in $(0,1)$ and admits a Puiseux expansion near $1$ as Lemma 4.4. Analogously to the proof of Lemma 4.7, we can show $\psi$ is even symmetric in direction $e$.
Remark B.1.
An interesting question is to extend the study in Proposition B.1 when the domain is non-convex. More precisely, in bounded simply connected domains with mirror symmetry, determine whether analytic steady flows with single-signed vorticity should enjoy mirror symmetry.